OpenStudy (anonymous):

Help please! (: Consider the equation: 2(x^2+y^2)^2 = 25(x^2-y^2) 1.) verify that the point P(3,-1) lies on the curve given by the equation 2.) Use implicit Differentiation to solve for dy/dx 3.) Find the equations of the tangent line and normal line at point P.

terenzreignz (terenzreignz):

let's do 1 first. At the left-hand side of your equation, you have \[\Large 2(x^2 + y^2)^2\] Now, substitute the point P(3,-1) by replacing x with 3 and y with -1 and tell me what you get...

OpenStudy (anonymous):

200?

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Okay, now try it for the right-hand side \[\Large 25(x^2 - y^2)\]

OpenStudy (anonymous):

200? @terenzreignz

terenzreignz (terenzreignz):

I hope you weren't guessing :) But yes, it is 200 on both sides when x = 3 and y = -1 Thus verifying that the point P(3,-1) is indeed on the curve :3

OpenStudy (anonymous):

haha no. I solved. so was that the answer for 1.? @terenzreignz

terenzreignz (terenzreignz):

The 'answer' to number 1 is to just replace x and y with 3 and -1 respectively and show that the equation holds.

OpenStudy (anonymous):

so it solving the two sides and having them equal each other? @terenzreignz

terenzreignz (terenzreignz):

yup :) Which they did, they both are 200...

OpenStudy (anonymous):

okay(: so #2? @terenzreignz

terenzreignz (terenzreignz):

There is no need to keep tagging me -.- Just call me Terence, okay? :P Now, you have to differentiate (both sides) of the equation... Can you do that?

OpenStudy (anonymous):

Sorry.... haha the right side would be 4(x^2+y^2)*(2x+2y) right? and the left side (100x-100y)?

terenzreignz (terenzreignz):

I think you've got your right and left sides switched... :P

OpenStudy (anonymous):

haha oh yeah. whoops.but did I derive them correctly?

terenzreignz (terenzreignz):

Nope :) Left one, almost... you just forgot that every time you differentiate the y bit, you affix a \(\large y'\) or a \[\Large \frac{dy}{dx}\] to it as per the chain rule...

OpenStudy (anonymous):

but my right one was correct? how do i derive the left?

terenzreignz (terenzreignz):

Left one is almost as you did, except... \[\Large 4(x^2+y^2)\left(2x + 2y\color{blue}{\frac{dy}{dx}}\right)\]

OpenStudy (anonymous):

i just needed to add dy/dx? so where so I go from there?

terenzreignz (terenzreignz):

Well, differentiate the right-side again, only this time, when you differentiate the y, include a dy/dx

OpenStudy (anonymous):

so it would be (100x-100y(dy/dx))?

terenzreignz (terenzreignz):

why 100? Last I checked, 2x25 is 50 :P

OpenStudy (anonymous):

omg. yes 25... so sorry my brain is not functioning well. Been doing math for 7 hrs now lol soo. (50x-50y(dy/dx))?

terenzreignz (terenzreignz):

Yup :) Let's use y' instead... it looks prettier that way, see :P \[\large 4(x^2+y^2)\left(2x + 2y \color{red}{y'}\right)= 50x +50y\color{red}{y'}\]

OpenStudy (anonymous):

oh okay... so is that the answer? or? lol

terenzreignz (terenzreignz):

No. Most unfortunately, the worst is yet to come... You have to solve for y'. What I mean is you have to get it alone on one side of the equation... good luck :P

OpenStudy (anonymous):

oh man... so both y' to one side?

terenzreignz (terenzreignz):

Yup... and solve...

OpenStudy (anonymous):

I'm stuck.. I dont know where to begin..

terenzreignz (terenzreignz):

Begin by bringing all the terms with y' to the right (or left) side and those without, to the other side.

OpenStudy (anonymous):

\[(4x^2+4y^2)(2x)-50x = 50y(dy/dx)-2y(dy/dx)\]

terenzreignz (terenzreignz):

You forgot something....something you didn't distribute...

OpenStudy (anonymous):

the 2x(4x^2+4y^2)?

OpenStudy (anonymous):

6x^2+2x(4y)^2-50x = 50y(dy/dx)-27(dy/dx)

terenzreignz (terenzreignz):

Hang on... and for goodness' sake, use y' for now, not that menacing looking dy/dx :D

OpenStudy (anonymous):

hah okay. \[6x^3+(2x)4y^2-50x = 50yy'-2yy'\]

OpenStudy (anonymous):

where do I go from here? hah

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

It was wrong... for one thing, 2x4 is not 6, it's 8. Also, you forgot that this part also has a (x^2 + y^2) \[\large 8x^3+(2x)4y^2-50x = 50y\color{red}{y'}-2y(x^2+y^2)\color{red}{y'}\]

terenzreignz (terenzreignz):

But other than that, everything's fine...

OpenStudy (anonymous):

Wow, I'm really bad at this. okay. soo \[8x^3+(2x)4y^2-50x = 50yy'- (2y)x^2+2y^3(y')\]

terenzreignz (terenzreignz):

Hold it there... let's start from here first... \[\large 8x^3+(2x)4y^2-50x = 50y\color{red}{y'}-2y(x^2+y^2)\color{red}{y'}\]

terenzreignz (terenzreignz):

And factor out the y' in the right-side \[\large 8x^3+(2x)4y^2-50x = \color{red}{y'}\left[50y-2y(x^2+y^2)\right]\]

OpenStudy (anonymous):

then divide right? to get y' on its own?

terenzreignz (terenzreignz):

Yes.... So you get this really ugly expression.. \[\Large \color{red}{y' }= \frac{8x^3 + 8xy^2-50x}{50y-2y(x^2+y^2)}\] Oh well... I suppose it could be worse :)

terenzreignz (terenzreignz):

And there's your answer :P

OpenStudy (anonymous):

wow. that was difficult lol I really appreciate the step by step process(:

terenzreignz (terenzreignz):

no problem

OpenStudy (anonymous):

So how do I go about finding the tangent and normal line at point P (3,-1)

terenzreignz (terenzreignz):

You just need slopes. to get the slope of the tangent line at the point P(3,-1), you... evaluate y' when x = 3 and y = -1 get to it, champ :)

OpenStudy (anonymous):

so just plug in x and y into y'?

terenzreignz (terenzreignz):

Yup. Have fun.

OpenStudy (anonymous):

-3 ?

OpenStudy (anonymous):

so tangent line would be. \[y=-3x-10\]

terenzreignz (terenzreignz):

ehh? y + 1 = -3(x - 3) again please...

OpenStudy (anonymous):

y+1=-3x-9

OpenStudy (anonymous):

then you minus 1 from y. which makes the -9 a -10?

terenzreignz (terenzreignz):

+9 not -9 sheesh what's negative times negative? :D

OpenStudy (anonymous):

i meant y+1=-3x+9 so y=-3x+8 ?

terenzreignz (terenzreignz):

Yes, much better. Now for the normal line, the slope is just the negative of the reciprocal of -3

OpenStudy (anonymous):

wait huh? negative of the reciprocal of -3? so 1/3?

terenzreignz (terenzreignz):

yup :) That slope, but passing through the point (3,-1) that's the normal line :P

OpenStudy (anonymous):

so y=1/3x-2 ?

terenzreignz (terenzreignz):

good. Well done :P

OpenStudy (anonymous):

Yay! (: thank you so much Terence! I really appreciate all the help!

terenzreignz (terenzreignz):

This is assuming you evaluated y' correctly... but I'm sure you did :P

OpenStudy (anonymous):

I hope so lol